16/17 (Feb) TEL Programme Monthly Meeting

TEL Monthly Programme meeting (for Feb 2017 project info)

21 March 2017, 1100-1200

Present: Anne-Marie Scott, Wesley Kerr, Jo Kinsley, Geir Granum, Colin Watt

Apologies: Adam Wadee

 

Updates on In Progress Projects

TEL031 - following onsite meetings with Civitas w/c 13/03 there is expectation of signoff of outstanding issues by end March; this includes signoff on data validation as applied to programmes explored and set of steps which can be repeated for new programmes. Expectation that there will be meetings with Civitas on site / on line end March / Early April to confirm and to look at courses tool.

TEL033 - as the majority of development work was expected to be done by SSP, and Director of LTW has been holding funds to transfer to SSP; Programme Owner has been awaiting steer on feasibility of doing work this session but has not had confirmation; Programme Owner would still like to explore Keith Brunton's CSV conversion tool which takes agreed exchange exports from Learn and puts into CSV format for upload to EUCLID; therefore agreement with Programme Owner is to suspend the project, hold 50 days in programme on balance row), and determine feasibility of CSV work to produce slicker tool to take extract from Learn that is EUCLID ready; to be reviewed at next programme meeting and agree scope change.

TEL034 - universe to be promoted to test this week for additional checks when Jenni returns; proposal to cover days for closure activity from programme (tbc).

TEL042 - Options appraisal for review w/c 27/03/17 and on track for decision and subsequent project closure mid-April.

TEL043 - work progressing on building blocks more quickly than expected; no concerns on budget at present.

TEL044 - discussed PM resource conflict; change of PM discussed with ISG Portfolio Manager. Programme Manager to confirm PM asap.

TEL045 - UoE theme to be done; meeting arranged to agree who is doing the work; discussed PM resource conflict; change of PM discussed with ISG Portfolio Manager. Programme Manager to confirm PM asap.

 

Updates on Not Started Projects

QMP upgrade (not started) As QMP have not provided a release date for v6 on-premise, the feeling is that this work is unlikely to be done in 16/17 however that if QMP were to release it May/Jun '17 then the project could be started with a view to deploying in semester 2 (17/18). In this case, a number of days could be declared as carryover for 17/18. This had been entered as a project in 17/18 3 year planning as contingency in case it was not possible to complete in 16/17 session (which is now increasingly likely); in meeting with ISG Portfolio Manager and TEL Programme Manager Jan 2017 the Head of Project Services had requested this be removed from the 17/18 plan and that if it was required as an in-year compliance project during that session then it would be funded. Programme Manager to request confirmation from Head of Project Services that this is the case, and guidance on planned carryover of days 16/17 to 17/18. Proposal is to change budget to 50 days for 16/17 project and reduce to with 50 days carry over to 17/18 (dependant on guidance from Head of Project Services on Annual Planning); there will also need to be a "go / no-go" decision date (May ?) at which point the call can be made on moving all the work into 17/18. Jo to speak with Mark Findlay on eta for plan / release schedule to assist this decision.  

 

AOB: 3 Year Annual Planning - Programme Owner requested update on communication of plan; Programme Manager to check there has been no comms missed.  

 

Actions:

Colin to confirm changes to TEL033 and request project suspension. (done)

Colin to organise meeting to discuss feasibility of work on Keith Brunton's CSV tool. (done)

Colin to confirm days for TEL034 completion and closure activity can be covered by TEL programme. (done) 

Colin to confirm change of PM for TEL044 Pebblepad upgrade. (done)

Colin to confirm change of PM for TEL045 Moodle upgrade. (in progress, tbc)

Colin to confirm steer from Head of Project Services on planned carryover of days to 17/18. (in progress)

Jo to check eta for QMP v6 release

Colin to check comms on 3 year annual planning. (done)